On a frictionless horizontal air table, puck A (with mass 0.254 kg ) is moving toward puck B (with mass 0.366 kg ), which is initially at rest. After the collision, puck A has velocity 0.119 m/s to the left, and puck B has velocity 0.650 m/s to the right.

1. What was the speed vAi of puck A before the collision?

2. Calculate ΔK, the change in the total kinetic energy of the system that occurs during the collision.
Answer #1

a)

final velocity of the puck A is given by

v2 = (m1 - m2) u1 / ( m1 + m2)

- 0.119* ( 0.366 + 0.254) = ( 0.254 - 0.366) u1

u1 = 0.65875 m/s, to the right

========

b) initial kE

kEi = 0.5* m1 u1^2 = 0.5* 0.254* 0.65875^2 = 0.0551 J

final KE

kEf = 0.5* ( 0.254* 0.119^2 + 0.366* 0.65^2) = 0.07911 J

change in kE = 0.024 J